What is the resistance of a 150m wire made of silver with cross sectional area of 0.60mm² ?


1
Expert's answer
2022-04-05T19:19:17-0400

The resistance of a wire

"R=\rho\frac{l}{A}"

"R=\rho\frac{l}{A}=1.59\times10^{-8}\times \frac{150}{0.60*10^{-6}}=4.0\:\Omega"